# Introducing Spotlight: A Novel Approach for Generating Captivating Key Information from Documents

### v1

@@ I.6 Cross-Domain Generalization of DPO
Our primary focus is on establishing a solid foundation for spotlight generation and demonstrating the efficacy of our proposed method within the chosen domains. Table 5 shows a considerable performance improvement of models when trained using SFT or DPO on a specific domain, but a question remained on the generalizability capability of the proposed models. We take the best-performing model (Llama3-8b DPO model) trained on one dataset and test it on the other dataset. An additional set of experiments in Table 29 shows the generalizability of the best performing model, DPO (L3-8b) from Table 5. indicating that the DPO-enhanced model robustly generalizes to unseen domains without significant degradation. This outcome reinforces the model’s consistency and effectiveness across varied datasets. The results show comparable outcome to in-domain dataset assessment as shown in Table 5.

### v3

@@ I.6 Cross-Domain Generalization of DPO
Our primary focus is on establishing a solid foundation for spotlight generation and demonstrating the efficacy of our proposed method within the chosen domains. Table 6 shows a considerable performance improvement of models when trained using SFT or DPO on a specific domain, but a question remained on the generalizability capability of the proposed models. We take the best-performing model (Llama3-8b DPO model) trained on one dataset and test it on the other dataset. An additional set of experiments in Table 30 shows the generalizability of the best performing model, DPO (L3-8b) from Table 6. indicating that the DPO‑enhanced model robustly generalizes to unseen domains without significant degradation. This outcome reinforces the model’s consistency and effectiveness across varied datasets. The results show comparable outcome to in-domain dataset assessment as shown in Table 6.